OpenStudy (anonymous):

Need solutions for the checkpoint questions in the slide show for Physics I : (fall 2010)

OpenStudy (experimentx):

could you point out particular problem?

OpenStudy (anonymous):

These 2 questions please

OpenStudy (experimentx):

I think you need to use the conservation of angular momentum for first question.

OpenStudy (anonymous):

We are expected to do the question from the concept of work just before this question they explained about dot product.

OpenStudy (experimentx):

what force do you have to apply to bring that rotating ball closer?

OpenStudy (experimentx):

it's centripetal force ... he centripetal force is given by \[ F = mv^2/r\] to bring it to R_1 distance from R_2 \[ \int_{R_2}^{F_1} F \cdot dr \] Now equate this thing with KE.

OpenStudy (anonymous):

\[mv ^{2}\ln (r2/r1)=mv ^{2}/2\] m, v2 cancels giving \[\ln (r _{2} div r _{1})=0.5\] so, it is independent of v; thus, v should remain constant am i correct experimentX

OpenStudy (experimentx):

hmm .. no mvr = J <-- this is constant. so v will be function of 'r'

OpenStudy (anonymous):

how did u arrive to this conclusion need a step by step guide

OpenStudy (experimentx):

hmm ... we know that angular momentum remains converged ... no matter what.

OpenStudy (anonymous):

angular momentum is mvr which should be cpnsta got it thanx